wiki:UC20210406

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 04/06/2021

Attending: Nancy Chanover (NMSU), Jamey Eriksen (APO), John Bally (CU), Adam Kowalski (CU), Mukremin Kilic (OU), Russet McMillan (APO), Kevin Schlaufman (JHU), Misty Bentz (GSU), Eric Bellm (UW), Denise Stephens (BYU), Anne Verbiscer (UVa), Joanne Hughes (Seattle U), Bill Ketzeback (APO), Ben Williams (UW), Sarah Tuttle (UW), Derek Buzasi (FGCU)


User feedback and comments from institutional representatives

Welcome to Adam Kowalski from CU, who will be taking over as CU's Users Committee rep and institutional scheduler in June!

  • NAPG - no report
  • Seattle - nothing to report
  • FGCU - nothing to report
  • NMSU - nothing to report from users; Moire is still working to understand an observed decrease in sensitivity in the far blue.
  • Washington - nothing to report
  • Oklahoma - Muk's class had a great virtual training session - special thanks to Candace and Russet for doing such a great job
  • UVa - nothing to report
  • Georgia State - nothing to report
  • JHU - users are happy with the new ThAr lamp for ARCES and the resulting wavelength calibrations
  • Wyoming - no report
  • Colorado - nothing to report; they also had a class training a few weeks ago
  • BYU - nothing to report

Telescope and Instruments Report

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 3/01/2021 - 4/01/2021

0) Overview

APO transitioned to a 33% staffing limit on 3/10/21, when Otero County moved from the Red to Yellow COVID tier. The State of New Mexico evaluates the tiered positions every two weeks on Wednesdays for each county in NM. Otero County remained Yellow on 3/24, and the next update will be on 4/7/21. Currently, Otero County has a 3.63% positivity rate and 13.30 cases per 100k.  For comparison, Dona Ana County (where NMSU is located), which is also in the Yellow tier, has a 4.06% positivity rate and 15.20 cases per 100k.

The data and numbers are available here: https://cv.nmhealth.org/public-health-orders-and-executive-orders/red-to-green/

We are still restricting all visitors to the site. Following the State of New Mexico and NMSU guidelines, there are travel restrictions for staff as well as visitors.

Weather for the month of March has been warming, with occasional wintery weather.

1) Telescope

The telescope is operational with no major issues this last month. 

2) Instruments

DIS: The scattered light in the red camera is starting to change but is still considered good. The blue camera is stable and is also still considered good.

TripleSpec: powered on, cold and operational.

Agile: powered on, cold and operational.

Echelle: powered on, cold and operational. The inter-order light ratio (IOL) for red and blue is stable and considered good. The ThAr calibration lamps were replaced, which increased the intensity of some of the fainter lines.

NICFPS: powered on, cold and operational. 

ARCTIC: powered on, cold, and operational. Issue with dome flats is being debugged and solutions are being worked on.  (“Fence” baffle on the mirror cover petal design is being manufactured at UVa machine shop.  Prototype testing hopefully in coming weeks.). The new rotation redesign for the diffuser have arrived on site and install will begin when staffing levels permit, currently planned for mid/late April install.

3) ARCSAT

ARCSAT is operational and has been used for observations this past month.   Off-axis guiding is available and recommended for exposures greater than 10 minutes.  Shelyak eShel commissioning is ongoing currently, as well as scheduled for more May 17-30. 


Additional telescope and instrument discussion

We learned today that KPNO had a successful remediation of asbestos in their aluminization chamber, and they are bringing it back online this month. This is relevant to us because once the facility returns to operational status we can schedule the realuminization of the 3.5m primary mirror, which in turn will drive our summer shutdown schedule. We will keep this group posted as we learn more; hopefully we will have more news by our next meeting.

Russet brought up a rapidly evolving issue regarding DIS. Historically we have seen the scattered light in DIS get worse in the summer with warmer temperatures. We are just now barely beginning to see hints of that, but there is also a readout difficulty that appears to be temperature related. We are not sure if it is associated with changes in the chip temperature or the temperature of the readout electronics. This year the chip temperature is running warmer than normal, and we may reach the threshold temperature above which we start seeing problems this month if our warm weather trend continues. If the problem is with the readout electronics it likely will not appear until June. Bill and Russet will continue to investigate this. Our recommendation for DIS users  (especially those observing in A half) is to ask the observing specialist if there are any updates, and/or to look for updates posted on the instrument news page (https://www.apo.nmsu.edu/arc35m/Engineering/index.html). The recommended way to deal with this is to acquire darks as part of one's observations and subtract off the unwanted signal in post-processing (normally darks are not needed for DIS).


KOSMOS Update

Sarah and team are in the process of closing up the dewar with the detector now installed. They will then begin the process of pumping it back down, which will take about a week, and will then move towards lab testing. They plan to do as much testing as possible in lab prior to sending the instrument down to APO for commissioning.


ARCSAT/SurveyCam Update

Users have been reporting a larger number of issues with SurveyCam in the past few months. The most significant of these is some streaking in the images that does not appear to flat-field out. The detector is old and getting it serviced is cost-prohibitive since the camera is no longer supported by the vendor, so we don't have a long-term solution to this problem. For some applications (e.g. astrometry) this is more of an annoyance than a show-stopper, but for others it could be more problematic. Other issues that have cropped up seem to be related to the camera/telescope control software and we continue to run diagnostics during the daytime (as time permits, which is challenging since we are still limited to 33% of our normal on-site personnel staffing levels) to see if we can get to the bottom of it. Thus, users should be made aware that SurveyCam is having these issues and should be viewed as usable in a shared-risk mode only. Everyone scheduled on ARCSAT during Q2 is aware of this situation, but this may impact instrument selection for proposers in future quarters.


ARCTIC Diffuser Rotation

The rotation stage for the ARCTIC diffuser is nearly complete. It will be integrated with the robotics box on the instrument later this week, then some software modifications will be needed to integrate it into TUI and the ICC. The engineering staff repaired the linear stage problem that was reported previously, which means that we should no longer have to turn the instrument upside down in order to move the diffuser in and out.


TUI on New Mac Processor

In response to concerns expressed about various (not APO-specific) software packages being incompatible with the new Mac M3 processor, we ran some tests using a new machine to ensure that TUI is still functional on computers with the new chip. It seems to be, but if users encounter any issues please report them to your observing specialist.


Q2 3.5m scheduling

There is still some open time available in the Q2 schedule (one short A half slot on April 30 and some more in May and June). Users are reminded to follow the standard procedure of emailing requests for this time to Ben, Russet, Nancy, and their institutional scheduler.


Q2 ARCSAT scheduling

There is no open time available in Q2.


3.5m Science Symposium

Nancy reopened the discussion of holding a virtual APO science symposium. This idea was discussed several months ago but at the time people felt that everyone was overloaded with online classes and general pandemic-induced stress so we elected to postpone the discussion to a later date. This time there was greater enthusiasm for the idea, because (a) it would be nice to have an opportunity to discuss science, (b) it would be a relatively low stakes forum through which students could present results from their research projects involving APO data, and (c) it would be useful to hear what people at other ARC institutions are working on so that we can explore options for cross-institutional collaborations (and potentially pooling of observing time). There was some discussion concerning when would be the ideal time for this meeting. Those on semester systems argued for late July since the fall semester activities begin ramping up in mid-August. Those on the quarter system argued for late September, which would be the end of their summer quarter. Nancy will consider date options and begin some early planning. We need to get the word out early so that people can reserve the dates and so that we can encourage broad participation within our departments.


ACTION ITEMS

Open action items from previous meetings:

  • Develop a survey to get feedback on the virtual training session - Nancy. Status: CLOSED. Survey is completed.

New action items from this meeting:

  • UC reps - notify DIS users of the potential for increased background and steps to check/remediate the issue.
  • UC reps - notify ARCSAT users of SurveyCam status.
  • Nancy - send completed training survey to Muk, Chip, Denise Stephens, Kevin France, and Kevin Schlaufman.
  • Nancy - finalize date(s) for virtual science symposium so we can announce it no later than next month.

All Other Business

None.


Next meeting

The next meeting will be on Tuesday May 4 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.